Sandhill Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $50 throughout the country to loyal alumni of over 2,100 schools. Sandhill’s variable costs are 40% of sales; fixed costs are $118,000 per month.

3c) Assume that variable costs increase to 45% of the current sales price and fixed costs increase by $15,000 per month. If Sandhill were to raise its sales price by 10% to cover these new costs, what would be the new annual breakeven point in sales dollars? (Round sales price to 2 decimal places, e.g. 52.75 and final answer to 0 decimal places, e.g. 5,275.)
Answer #1

Answer

  Statement Showing The Break- Even Sales at the Current Sale Price

Description Amount($)
Sale Value Per Unit 50
Less: Variable Cost @ 40% of Sales Value 20
Contribution Per Unit 30
Profit Volumn Ratio = (Contribution / Sales Value) 0.6
Fixed Cost for the Period $ 118000
Break- Even Sales 196667

Now the revised Sales Price Per Unit, Variable Cost and Fixed Cost will be -

Revised Sales Price = 50 + ( 50 x 10/100) = $ 55/ Unit

Revised Variable Cost Per Unit = ( 50x 45/100) = $ 22.50 / Unit

Revised Fixed Cost Per Unit = ( 118000 + 15000) = $ 133000.

Statement Showing the new Break- Even Point Sales

Description Amount ($)
New Sales Price Per Unit $55.00
Less: Variable Cost Per Unit $22.50
Contribution Per Unit $32.50
New Profit Volumn Ratio ( 32.50/55.00) 0.59
New Fixed Cost for the Period $133000
New Break-Even Sales ( New Fixed Cost / New P/V Ratio) $225424